Jack's stride is 64 cm long and Jill's is 56 cm long. If they start together, how far have they walked when they are next in step?
Answer by mananth(16946)

Jack's stride is 64 cm long
Jill's is 56 cm long.
..
Let them be in step after x strides
time taken by jack for x strides = x/64
..
time taken by Jill for x strides = x/56
..
x/56-x/64=0
64x-56x= 64*56
8x = 64*56
x= 64*56/8
x= 8*56
x=448
..
It is just like taking LCM of 64 & 56